# What is emergence, after all?

### v3

@@ Abstract
This manuscript was compiled on October 7, 2025
@@
Emergence occurs where a whole exhibits properties absent from its individual parts—an idea that traces back to Aristotle’s dictum that “the whole is something besides the parts” (1). In science, the term refers to patterns, structures, or behaviors that are not straightforwardly deducible from components in isolation (2, 3, 4, 5, 6, 7), yet are often captured by coarse, predictive descriptions that do not track every micro-detail. You can see emergence in the wild—flocks wheeling as one, fireflies locking into rhythm, insect societies coordinating without commanders (8)—and through computer simulations, where minimal rules produce striking order: Reynolds’ “boids” (9), cellular automata (10), and the Game-of-Life glider (11).
@@
Philosophical roots of the concept trace back to Mill’s distinction between additive “mechanical” effects and qualitatively new “chemical” ones (12). Lewes later coined the term “emergent” for such cases (13), and early British emergentists (14, 15, 16, 17) developed it into a layered view of nature in which higher-level laws complement, and sometimes influence, lower-level processes. This tradition faded as reductionism gained ground, driven by the intertheoretic reduction models (18, 19, 20). The debate revived with arguments for multiple realizability (21, 22, 23) and Anderson’s claim that “more is different” (24), followed by work on supervenience and causal autonomy (25). In recent years, the term has evolved from describing specific transitions, such as the “emergence of metabolism” (26), to becoming a popular buzzword, gaining more usage in scientific and non-scientific discourses (27, 24). Fig. 1 shows this trend.
@@
Here, we review operational perspectives on emergence—consistent with physicalism and grounded in effective theories and information-theoretic principles—rather than non-mainstream viewpoints (28, 29, 30, 31, 32, 33, 34). In the sections that follow, we assemble the concepts required for a working account of emergence across physics, biology, and the social sciences. We formalize emergence through lossy compression that preserves the information relevant for prediction and control, and we motivate Effective Theories as the mechanism that secures autonomy and multiple realizability. We then treat quantitatively secure testbeds—critical phenomena and phase transitions—to show how order parameters, response functions, scaling, and data collapse locate the onset of emergence and render it mechanistic and testable. Next, we use dualities to show that distinct microdescriptions can yield the same macrophysics, grounding autonomy without extra causes. Finally, we distinguish between explanatory and ontological reduction: reduction secures cross-scale consistency, while effective theories deliver explanations and control. Sequenced this way, the paper justifies an economy-of-description perspective: we can explain how many emergent phenomena happen—even if not yet all—and none require supernatural causes.
@@
Knowledge unfolds across successive levels of abstraction—from raw experience to language and higher-order concepts (36). Confusion between these levels distorts understanding. Emergence is everywhere, and the surprise it evokes often reflects where we choose to stop asking more questions (37, 25). What we call “emergent” typically depends on the limits of our knowledge, tools, or perspective—in other words, on epistemology (38, 39). Mixing red and green paint yields a dull brown—basic chemistry and optics explain it well enough for most, and few would call the result emergent. Yet a physicist who tracks the same blend down to quantum‐mechanical interactions would hardly find the task trivial.
Explanation is always a matter of perspective: an economist analyzing a financial crisis need not invoke atoms or molecules but instead works at the social scale relevant to the questions at hand. Every phenomenon can be analyzed at multiple descriptive levels, and what feels emergent at one level may be obvious at another (38). Even so, emergent behaviors are not merely in the eye of the beholder—they can display objective, quantifiable signatures such as information flow or causal strength (40, 41). We can make emergence more rigorous by formalizing it through coarse-graining maps that discard detail yet preserve prediction.
@@
Emergence is present when there exists a many-to-one map from a micro-level theory (more fundamental, detailed, or lower-level) to a macro-level theory (higher-level), such that the macro description remains predictive even after discarding most of the microscopic detail (42). As Fig. 2(a,b) presents, a map is a selective, structured, purpose-driven abstraction—useful precisely because it is not the whole territory it depicts (36, 43). Every map omits, distorts, or abstracts; acknowledging those sacrifices is part of scientific hygiene (43).
@@
where $N^{\prime}<N$, and the coarse-grained observables $\{X_{j}\}$ may be defined on a different (typically coarser) time grid $\{t^{\prime}_{\ell}\}$. Therefore, $\mathcal{F}$ is a lossy data compression where only the information needed for prediction survives across scales (45, 29). Fig. 2 pictures such maps. This viewpoint offers a substrate-independent vocabulary for linking small-scale descriptions to large-scale laws. In high-energy physics, the information relevant for prediction determines which effective theory you use at each scale (46). In brains and societies, the same approach helps identify mesoscales that are maximally compressed yet still predictive (47, 48, 49, 50, 51). Panels (a) and (b) show two different sets of coarse-graining where they average out the observables. Panel (e) illustrates how, by compressing data through the right process, we arrive at a simpler theory. The map $\mathcal{F}$, however, need not be simple, trivial, local, or unique.
@@
The first step toward understanding how emergence occurs is to identify when and under what conditions it arises, as well as its immediate consequences. Useful large-scale descriptions may need entirely different degrees of freedom than the ones appearing in the small-scale theory, and they can realize symmetries in new ways. Choosing the right variables is a scientific discovery step, not a fixed recipe. The emergence of life, mind, or social structure appears puzzling not because they conflict with physicalism, but because they represent a qualitatively more complex class of emergent phenomena (52). Therefore, there is no universal recipe for identifying the “right” macroscopic variables. The key practical questions here are whether there exists a finite set of macroscopic parameters that can adequately describe the phenomena of interest within a specified tolerance, and if so, which mapping $\mathcal{F}^{\star}$ minimizes the chosen distortion measure while achieving that macro-level description. The outcome is a systematic and metric-aware approach to transition from micro-level dynamics to macro-level dynamics.
@@
Figure 2: Coarse-graining, Universality, and RG-flow. (a) Temperature as a coarse-grained variable: under local equilibrium, averaging kinetic energies defines temperature, $\mathcal{F}:(v_{1},\ldots,v_{N})\!\mapsto\!T$. (b) Local coarse-graining by $2{\times}2$ blocks with majority-vote, $S_{I}=\operatorname{sgn}(s_{i}+s_{j}+s_{k}+s_{l})$ for $s_{\ell}\in\{\pm 1\}$; iterating the map flows toward a uniform block-spin configuration. (c) Liquid–gas criticality shares the Ising universality class. The control parameter is the reduced temperature $\tau=(T-T_{c})/T_{c}$. On a lattice, the lattice-gas mapping $n_{i}\!\in\!\{0,1\}\mapsto\sigma_{i}=2n_{i}-1$ (occupied $\uparrow$, empty $\downarrow$) makes the density deviation $\Delta\rho=\rho-\rho_{c}$ directly analogous to the magnetization $m$. (d) Increasing temperature drives a continuous transition from an ordered phase to a disordered phase, and the transition point we see scaling behaviors (28, 53). (e) Coarse-graining compresses high-resolution data and induces a renormalization-group (RG) flow in theory space. When the coarse-grained description is related to the microscopic one by parameter redefinitions, the model is said to be renormalizable. A macro-law screens off micro-details but remains autonomous, meaning that macro variables aren’t just convenient summaries but “real enough” to be the subject of genuine laws and explanations—even if many different microstates realize the same macrostate.
@@
Each scale brings qualitatively new behaviors that demand their own inquiry. So, reducing a phenomenon to fundamental laws does not mean we can reconstruct everything from them. Reductionism is not wrong—it is just insufficient for understanding the universe. Even a complete microscopic “Theory of Everything” would leave many of the essential rules governing higher-level systems untouched (55). Macroscopic behavior is determined by emergent parameters that are largely insensitive to the fine details of the underlying microscopic governing equation (56). Unifying quantum mechanics and gravity would undoubtedly be a milestone in Fundamental Physics, but it would not immediately explain why financial markets crash. These are collective phenomena with different ontologies, which cannot be derived by reduction alone (55).
@@ A New Ontology
Classical Mechanics itself may emerge from a quantum theory, which is, in turn, ontologically different (61). In Classical Mechanics, a particle is represented by a point in phase space with well-defined position and momentum. In Quantum Mechanics, the ontology shifts: a particle is represented by a vector in Hilbert space, and is described by a wavefunction (62). Position and momentum are no longer coordinates but observables. The uncertainty principle ensures that no wavefunction can define both position and momentum with perfect precision. Thus, quantum mechanics offers no definitive answer to the classical question of “where” a particle is, or “how fast” it is moving (62). Nevertheless, the classic macroscopic theory remains applicable in its domain, and it is consistent with the underlying quantum description through the processes of decoherence and measurement, which effectively produce the appearance of wavefunction collapse (63).
@@
In the large-size limit, Thermodynamics is not merely an approximation to Statistical Mechanics—it is an effective theory (65), one that captures the essential behavior of macroscopic systems using a few key variables at the right scale (66, 67). Effective theories isolate what matters at a given scale and discard what does not (45). They are not shortcuts; they are self-contained, predictive, and often universal descriptions of emergent levels of reality. That is why, for specific questions, instead of tracking the detailed positions and velocities of every particle, we use variables like temperature, pressure, and entropy with their own simple and powerful macro-level laws (68, 69). A mean-field theory is the zeroth-order effective description (70). It is invaluable as a first cut, exact when fluctuations vanish, but insufficient whenever correlations hold the key to collective behavior (71).
Not all effective theories are created equal. In some systems—particularly those involving biological, cognitive, or social dynamics—the coarse-grained variables do not merely summarize the current state of the underlying components; they also retain memory, storing information about past configurations. Such memory and nonlocality are not exclusive to living or social systems: in physics, integrating out fast fields generically produces spatially nonlocal interactions and temporal memory kernels (72, 73, 74). Even gravity has been cast in explicitly nonlocal, memory-like form (75, 76, 77). This historical dependence alters the structure of the effective theory, introducing higher-order terms that reflect feedback, path dependence, and self-reference (42). Such systems are still governed by physical laws, but their dynamics can become computationally intractable, meaning that no tractable microscopic derivation will fully recover the macroscopic behavior (52).
@@
A clear and intuitive example of how emergence happens is magnetization. In a fridge magnet, the collective alignment of billions of electron spins (little magnets) produces a macroscopic magnetic property, denoted $m(T)$, which appears only when the system is below a critical temperature $T_{c}$. For a fridge magnet, it is around $450^{\circ}$C. Heating the material beyond this point destroys the alignment, and as Fig. 2d shows, the magnetic property disappears in a continuous phase transition (28). In other words, heating injects fluctuations that destroy spin alignment, driving the system from a magnetically ordered phase ($m>0$) to a disordered paramagnetic phase ($m=0$). A single governing equation (Ising Hamiltonian) describes both magnetized and non-magnetized states (28).
@@
At the critical point $T=T_{c}$, the spin-flip symmetry of the Hamiltonian remains intact, but the actual state chooses a specific direction, resulting in $m\neq 0$ (70, 78). This symmetry breaking defines an order parameter that compresses the full $10^{23}$-spin microstate into a single coarse-grained vector whose dynamics—like domain walls or spin waves—obey new effective laws (79, 80, 24). At $T_{c}$, key properties of the system—such as correlation length and relaxation time—diverge (or peak in finite systems), with finite-size scaling and data collapse as empirical checks (81, 82, 28). These divergences, characterized by critical exponents, serve as clear markers of the onset of emergence.
@@
More interestingly, different systems composed of distinct elements can exhibit the same critical behavior (83). That is, systems with very different microstructures can fall into the same universality class and be described by the same macroscopic theory. This substrate-independence also means that, from the perspective of the emergent behavior, one cannot tell which microscopic system produced it. When very different microscopic models display the same large-scale behavior, even exceptionally precise measurements of large-scale behavior usually cannot identify a unique microscopic theory without extra assumptions (e.g., symmetry content or locality) (84, 81).
@@
The liquid–gas transition belongs to the same universality class as the ferromagnetic transition (53). Fig. 2c illustrates the technique of translating these two transitions into one another. Universality classes, therefore, organize not just critical exponents but also the very symmetry content that survives at macroscopic scales, reinforcing why broken symmetry is one maker of emergence (24, 54). In this sense, “more is different” (24): once many degrees of freedom lock together, they generate effective laws—and sometimes entirely new effective symmetries—that are absent from, yet fully compatible with, the underlying dynamics.
@@
It is worth mentioning that information-theoretic diagnostics similarly pinpoint the onset of emergence: mutual information rises near criticality, transfer entropy reveals increasingly predictive interactions, and persistent mutual information isolates long-lived choices beyond short-time correlations (45, 49, 50). Causal tests complement these signals: Effective Information can increase under coarse-graining—so a well-chosen macro description exerts greater causal constraint than the micro—and partial information decomposition exposes downward causation and causal decoupling directly in multivariate data (85, 86).
@@
Technically, critical universality manifests only within the critical region of a continuous phase transition: as control parameters approach the transition, the renormalization-group (RG) flow is drawn toward a scale-invariant fixed point—attractive along irrelevant directions and unstable along relevant ones—so macroscopic behavior becomes largely independent of microscopic details (84, 87, 88). This picture applies to both equilibrium and nonequilibrium continuous transitions (89, 90); nothing forbids far-from-equilibrium universality (91), though for certain many-body Hamiltonians even basic features (spectral gap, phase diagram, etc.) are undecidable in principle despite a computable RG flow (92).
@@
A canonical non-equilibrium exemplar is the class of absorbing-state transitions (93, 94, 95): dynamics are irreversible, the system can enter configurations it cannot escape, and the order parameter (e.g., the active-site density in the contact process) vanishes continuously at a critical control parameter with power-law correlations at criticality (89, 90). Under broad conditions, these phase transitions fall into the directed percolation (DP) universality class, often regarded as the non-equilibrium analogue of the Ising paradigm we mentioned before (89, 90). Neural-population recordings, deep neural networks, and insect swarms also exhibit RG-style scaling in nonequilibrium settings (96, 97, 98, 99, 100, 101). What truly matters for RG are locality, symmetry, and conservation constraints, and dimensionality (89, 102). That said, beyond a few well-understood universality classes, we still lack a clear, well-founded theoretical framework—comparable in scope to the renormalization group—for systematically organizing and classifying non-equilibrium phases (89).
@@
One must take care to remember that criticality is neither the only route to emergence nor trivial to establish empirically (103). For example, it is not straightforward to extend the statistical mechanics framework of critical phenomena to describe abrupt qualitative changes in dynamical systems. Likewise, the emergence of life from chemical interactions does not neatly correspond to a phase transition in the physical sense, despite claims to the contrary by some physicists (104). Intricate phenomena such as consciousness or thought could, in principle, arise from entirely different physical substrates—biological neurons or silicon circuits alike (105, 106)—provided there is sufficient evidence that they result from well-defined critical phenomena (107). Across biological systems, current evidence favors near-criticality over exact self-tuned criticality (108, 109, 110). In neuroscience in particular, recent surveys argue that a distance-to-criticality perspective is more realistic and useful—for cognition and pathology—than the claim that the brain sits precisely at criticality; both edge-of-chaos and avalanche-like regimes can appear, often state-dependently (108, 109, 110). Therefore, “critical-looking” statistics (power laws, avalanches) are insufficient on their own (111, 96, 112, 113, 114) and broken symmetries underpin not all emergent behaviors (115, 116). Accordingly, strong claims of criticality, let alone emergence in living systems, require careful analysis (82, 117, 28, 89).
@@
Duality is an equivalence (often infrared-exact (118, 119)) between two seemingly different theories: a one-to-one “dictionary” that maps observables, states, and parameters so that all physical predictions coincide. Dualities reveal that theories with ontologically distinct variables—such as bosons versus fermions, or gravity versus no gravity—can describe the same underlying physics (118). Formally, a duality establishes a structural correspondence between the measurable quantities of two theories (a *-isomorphism) such that every observable, expectation value, and response function matches once external conditions—sometimes called “sources”—are translated accordingly. This agreement may hold up to local adjustments, like contact terms or counterterms, reflecting different renormalization schemes (118).
In 1+11{+}1 dimensions, the bosonic sine–Gordon model is exactly dual to the fermionic Thirring model, exchanging solitonic waves for interacting particles and relating their couplings nonlinearly (120). In the Anti–de Sitter/conformal field theory correspondence, a conformal field theory without gravity on a $d$-dimensional boundary is holographically equivalent to a (d+1)(d{+}1)-dimensional bulk theory with gravity (46). Classic examples from statistical and gauge theories—such as high–low temperature mappings, electric–magnetic duality, and particle–vortex correspondence—further illustrate how distinct microscopic descriptions can encode the same macroscopic behavior (121, 122, 123, 124)
@@
Moving beyond physics, we consider how large-scale patterns emerge from local interactions in social and informational systems. Emergent phenomena possess causal power. In both physical and social worlds, entities like society and culture act as real forces shaping outcomes. Even in systems designed without formal rules—such as decentralized communities like Wikipedia or Stack Exchange—structures, hierarchies, and power distributions inevitably emerge, and the resulting social laws profoundly influence the world.

### v4

@@ Abstract
This manuscript was compiled on January 9, 2026
@@
Emergence occurs where a whole exhibits properties absent from its individual parts—an idea that traces back to Aristotle’s dictum that “the whole is something besides the parts” (1). In science, the term refers to patterns, structures, or behaviors that are not straightforwardly deducible from components in isolation (2, 3, 4, 5, 6, 7), yet are often captured by coarse, predictive descriptions that do not track every micro-detail. Emergence can be observed in the world around us: polarisation in public discourse (8), flocks of birds moving in unison, fireflies synchronizing their flashes, and insect societies coordinating their activities without central command (9). It can also be studied in computer simulations, where simple local rules generate complex collective behavior—for example, Reynolds’ “boids” (10), cellular automata (11), and gliders in Conway’s Game of Life (12).
@@
Philosophical roots of the concept trace back to Mill’s distinction between additive “mechanical” effects and qualitatively new “chemical” ones (13). Lewes later coined the term “emergent” for such cases (14), and early British emergentists (15, 16, 17, 18) developed it into a layered view of nature in which higher-level laws complement, and sometimes influence, lower-level processes. This tradition faded as reductionism gained ground, driven by the intertheoretic reduction models (19, 20, 21). The debate revived with arguments for multiple realizability (22, 23, 24) and Anderson’s claim that “more is different” (25), followed by work on supervenience and causal autonomy (26). In recent years, the term has evolved from describing specific transitions, such as the “emergence of metabolism” (27), to a popular buzzword, gaining greater use in scientific and non-scientific discourses (28, 25). Fig. 1 shows this trend.
@@
Here, we review operational perspectives on emergence—consistent with physicalism and grounded in effective theories and information-theoretic principles (29, 30, 31, 32), whose thermodynamic meaning in conventional statistical mechanics is also well established (31, 33, 34). However, there is also a rich landscape of generalized non-additive entropies and nonextensive statistical mechanics (35, 36, 37), which provide alternative ways to describe emergence. We encourage interested readers to consult this literature for complementary perspectives. In the sections that follow, we assemble the concepts required for a working account of emergence across physics, biology, and the social sciences. We formalize emergence through lossy compression that preserves the information relevant for prediction and control, and we motivate Effective Theories as the mechanism that secures autonomy and multiple realizability. We then treat quantitatively secure testbeds—critical phenomena and phase transitions—to show how order parameters, response functions, scaling, and data collapse locate the onset of emergence and render it mechanistic and testable. Next, we use dualities to show that distinct microdescriptions can yield the same macrophysics, grounding autonomy without extra causes. Finally, we distinguish between explanatory and ontological reduction: reduction secures cross-scale consistency, while effective theories deliver explanations and control. Sequenced this way, the paper justifies an economy-of-description perspective: we can explain how many emergent phenomena happen—even if not yet all—and none require supernatural causes.
@@
Knowledge unfolds across successive levels of abstraction—from raw experience to language and higher-order concepts (39). Confusion between these levels distorts understanding. Emergence is everywhere, and the surprise it evokes often reflects where we choose to stop asking more questions (40, 26). What we call “emergent” typically depends on the limits of our knowledge, tools, or perspective—in other words, on epistemology (41, 42). Mixing red and green paint yields a dull brown—basic chemistry and optics explain it well enough for most, and few would call the result emergent. Yet a physicist who tracks the same blend down to quantum‐mechanical interactions would hardly find the task trivial.
Explanation is always a matter of perspective: an economist analyzing a financial crisis need not invoke atoms or molecules but instead works at the social scale relevant to the questions at hand. Every phenomenon can be analyzed at multiple descriptive levels, and what feels emergent at one level may be obvious at another (41). Even so, emergent behaviors are not merely in the eye of the beholder—they can display objective, quantifiable signatures such as information flow or causal strength (43, 44). We can make emergence more rigorous by formalizing it through coarse-graining maps that discard detail while preserving predictive power.
@@
Emergence is present when there exists a many-to-one map from a micro-level theory (more fundamental, detailed, or lower-level) to a macro-level theory (higher-level), such that the macro description remains predictive even after discarding most of the microscopic detail (45). As Fig. 2(a,b) presents, a map is a selective, structured, purpose-driven abstraction—useful precisely because it is not the whole territory it depicts (39, 46). Every map omits, distorts, or abstracts; acknowledging those sacrifices is part of scientific hygiene (46).
@@
where $N^{\prime}<N$, and the coarse-grained observables $\{X_{j}\}$ may be defined on a different (typically coarser) time grid $\{t^{\prime}_{\ell}\}$. Therefore, $\mathcal{F}$ is a lossy data compression where only the information needed for prediction survives across scales (32, 30). Fig. 2 pictures such maps. This viewpoint offers a substrate-independent vocabulary for linking small-scale descriptions to large-scale laws. In high-energy physics, the information relevant for prediction determines which effective theory you use at each scale (48). In brains and societies, the same approach helps identify mesoscales that are maximally compressed yet still predictive (49, 50, 51, 52, 53). Panels (a) and (b) show two different sets of coarse-graining where they average out the observables. Panel (e) illustrates how, by compressing data through the right process, we arrive at a simpler theory. The map $\mathcal{F}$, however, need not be simple, trivial, local, or unique.
@@
The first step toward understanding how emergence occurs is to identify when and under what conditions it arises, as well as its immediate consequences. Useful large-scale descriptions may need entirely different degrees of freedom than the ones appearing in the small-scale theory, and they can realize symmetries in new ways. Choosing the right variables is a scientific discovery step, not a fixed recipe. The emergence of life, mind, or social structure appears puzzling not because they conflict with physicalism, but because they represent a qualitatively more complex class of emergent phenomena (54). Therefore, there is no universal recipe for identifying the “right” macroscopic variables. The key practical questions here are whether there exists a finite set of macroscopic parameters that can adequately describe the phenomena of interest within a specified tolerance, and if so, which mapping $\mathcal{F}^{\star}$ minimizes the chosen distortion measure while achieving that macro-level description. The outcome is a systematic, metric-aware approach to transitioning from micro-level to macro-level dynamics.
@@
Figure 2: Coarse-graining, Universality, and RG-flow. (a) Temperature as a coarse-grained variable: under local equilibrium, averaging kinetic energies defines temperature, $\mathcal{F}:(v_{1},\ldots,v_{N})\!\mapsto\!T$. (b) Local coarse-graining by $2{\times}2$ blocks with majority-vote, $S_{I}=\operatorname{sgn}(s_{i}+s_{j}+s_{k}+s_{l})$ for $s_{\ell}\in\{\pm 1\}$; iterating the map flows toward a uniform block-spin configuration. (c) Liquid–gas criticality shares the Ising universality class. The control parameter is the reduced temperature $\tau=(T-T_{c})/T_{c}$. On a lattice, the lattice-gas mapping $n_{i}\!\in\!\{0,1\}\mapsto\sigma_{i}=2n_{i}-1$ (occupied $\bullet$, empty $\circ$) makes the density difference $\Delta\rho=\rho-\rho_{c}$ directly analogous to the magnetization $m$. (d) Increasing temperature drives a continuous transition from an ordered phase to a disordered phase, and the transition point we see scaling behaviors (29, 55). (e) Coarse-graining compresses high-resolution data and induces a renormalization-group (RG) flow in theory space. When the coarse-grained description is related to the microscopic one by parameter redefinitions, the model is said to be renormalizable. A macro-law screens off micro-details but remains autonomous, meaning that macro variables aren’t just convenient summaries but “real enough” to be the subject of genuine laws and explanations—even if many different microstates realize the same macrostate.
@@
Each scale brings qualitatively new behaviors that demand their own inquiry. So, reducing a phenomenon to fundamental laws does not mean we can reconstruct everything from them. Reductionism is not wrong—it is just insufficient for understanding the universe (57). Even a complete microscopic “Theory of Everything” would leave many of the essential rules governing higher-level systems untouched (58). Macroscopic behavior is determined by emergent parameters that are largely insensitive to the fine details of the underlying microscopic governing equation (59). Unifying quantum mechanics and gravity would undoubtedly be a milestone in Fundamental Physics, but it would not immediately explain why financial markets crash. These are collective phenomena with different ontologies, which cannot be derived by reduction alone (58).
@@ A New Ontology
Classical Mechanics itself may emerge from a quantum theory, which is, in turn, ontologically different (64). In Classical Mechanics, a particle is represented by a point in phase space with well-defined position and momentum. In Quantum Mechanics, the ontology shifts: a particle is represented by a vector in Hilbert space, and is described by a wavefunction (65). Position and momentum are no longer coordinates but observables. The uncertainty principle ensures that no wavefunction can define both position and momentum with perfect precision. Thus, quantum mechanics offers no definitive answer to the classical question of “where” a particle is, or “how fast” it is moving (65). Nevertheless, the classic macroscopic theory remains applicable in its domain and is consistent with the underlying quantum description through decoherence and measurement, which effectively produce the appearance of wavefunction collapse (66).
@@
In the large-size limit, Thermodynamics is not merely an approximation to Statistical Mechanics—it is an effective theory (68), one that captures the essential behavior of macroscopic systems using a few key variables at the right scale (69, 70). Effective theories isolate what matters at a given scale and discard what does not (32). They are not shortcuts; they are self-contained, predictive, and often universal descriptions of emergent levels of reality. That is why, for specific questions, instead of tracking the detailed positions and velocities of every particle, we use variables like temperature, pressure, and entropy with their own simple and powerful macro-level laws (71, 72). A mean-field theory is the zeroth-order effective description (73). It is invaluable as a first cut, exact when fluctuations vanish, but insufficient whenever correlations hold the key to collective behavior (74).
Not all effective theories are created equal. In some systems—particularly those involving biological, cognitive, or social dynamics—the coarse-grained variables do not merely summarize the current state of the underlying components; they also retain memory, storing information about past configurations. Such memory and nonlocality are not exclusive to living or social systems: in physics, integrating out fast fields generically produces spatially nonlocal interactions and temporal memory kernels (75, 76, 77). Even gravity has been cast in explicitly nonlocal, memory-like form (78, 79, 80). This historical dependence alters the structure of the effective theory, introducing higher-order terms that reflect feedback, path dependence, and self-reference (45). Such systems are still governed by physical laws, but their dynamics can become computationally intractable, meaning that no tractable microscopic derivation will fully recover the macroscopic behavior (54).
@@
A clear, intuitive example of emergence is magnetization. In a fridge magnet, the collective alignment of billions of electron spins (little magnets) produces a macroscopic magnetic property, denoted $m(T)$, which appears only when the system is below a critical temperature $T_{c}$. For a fridge magnet, it is around $450^{\circ}$C. Heating the material beyond this point destroys the alignment, and as Fig. 2d shows, the magnetic property disappears in a continuous phase transition (29). In other words, heating injects fluctuations that destroy spin alignment, driving the system from a magnetically ordered phase ($m>0$) to a disordered paramagnetic phase ($m=0$). A single governing equation (Ising Hamiltonian) describes both magnetized and non-magnetized states (29).
@@
At the critical point $T=T_{c}$, the spin-flip symmetry of the Hamiltonian remains intact, but the actual state chooses a specific direction, resulting in $m\neq 0$ (73, 81). This symmetry breaking defines an order parameter that compresses the full $10^{23}$-spin microstate into a single coarse-grained vector whose dynamics—like domain walls or spin waves—obey new effective laws (82, 83, 25). At $T_{c}$, key properties of the system—such as correlation length and relaxation time—diverge (or peak in finite systems), with finite-size scaling and data collapse as empirical checks (84, 85, 29). These divergences, characterized by critical exponents, serve as clear markers of the onset of emergence.
@@
More interestingly, different systems composed of distinct elements can exhibit the same critical behavior (86). That is, systems with very different microstructures can fall into the same universality class and be described by the same macroscopic theory. This substrate-independence also means that, from the perspective of the emergent behavior, one cannot tell which microscopic system produced it. When very different microscopic models display the same large-scale behavior, even exceptionally precise measurements of large-scale behavior usually cannot identify a unique microscopic theory without extra assumptions (e.g., symmetry content or locality) (87, 84).
@@
The liquid–gas transition belongs to the same universality class as the ferromagnetic transition (55). Fig. 2c illustrates the technique of translating these two transitions into one another. Universality classes, therefore, organize not just critical exponents but also the very symmetry content that survives at macroscopic scales, reinforcing why broken symmetry is one maker of emergence (25, 56). In this sense, “more is different” (25): once many degrees of freedom lock together, they generate effective laws—and sometimes entirely new effective symmetries—that are absent from, yet fully compatible with, the underlying dynamics.
@@
It is worth mentioning that information-theoretic diagnostics similarly pinpoint the onset of emergence: mutual information rises near criticality, transfer entropy reveals increasingly predictive interactions, and persistent mutual information isolates long-lived choices beyond short-time correlations (32, 51, 52). Causal tests complement these signals: Effective Information can increase under coarse-graining—so a well-chosen macro description exerts greater causal constraint than the micro—and partial information decomposition exposes downward causation and causal decoupling directly in multivariate data (88, 89).
@@
Technically, critical universality manifests only within the critical region of a continuous phase transition: as control parameters approach the transition, the renormalization-group (RG) flow is drawn toward a scale-invariant fixed point—attractive along irrelevant directions and unstable along relevant ones—so macroscopic behavior becomes largely independent of microscopic details (87, 90, 91). This picture applies to both equilibrium and non-equilibrium continuous transitions (92, 93); nothing forbids far-from-equilibrium universality (94), though for certain many-body Hamiltonians even basic features (spectral gap, phase diagram, etc.) are undecidable in principle despite a computable RG flow (95).
@@
A canonical non-equilibrium exemplar is the class of absorbing-state transitions (96, 97, 98): dynamics are irreversible, the system can enter configurations it cannot escape, and the order parameter (e.g., the active-site density in the contact process) vanishes continuously at a critical control parameter with power-law correlations at criticality (92, 93). Under broad conditions, these phase transitions fall into the directed percolation (DP) universality class, often regarded as the non-equilibrium analogue of the Ising paradigm we mentioned before (92, 93). Neural-population recordings, deep neural networks, and insect swarms also exhibit RG-style scaling in nonequilibrium settings (99, 100, 101, 102, 103, 104). What truly matters for RG are locality, symmetry, and conservation constraints, and dimensionality (92, 105). That said, beyond a few well-understood universality classes, we still lack a clear, well-founded theoretical framework—comparable in scope to the renormalization group—for systematically organizing and classifying non-equilibrium phases (92).
@@
One must take care to remember that criticality is neither the only route to emergence nor trivial to establish empirically (106). For example, it is not straightforward to extend the statistical mechanics framework of critical phenomena to describe abrupt qualitative changes in dynamical systems. Likewise, the emergence of life from chemical interactions does not neatly correspond to a phase transition in the physical sense, despite claims to the contrary by some physicists (107). Intricate phenomena such as consciousness or thought could, in principle, arise from entirely different physical substrates—biological neurons or silicon circuits alike (108, 109)—provided there is sufficient evidence that they result from well-defined critical phenomena (110). Across biological systems, current evidence favors near-criticality over exact self-tuned criticality (111, 112, 113). In neuroscience in particular, recent surveys argue that a distance-to-criticality perspective is more realistic and useful—for cognition and pathology—than the claim that the brain sits precisely at criticality; both edge-of-chaos and avalanche-like regimes can appear, often state-dependently (111, 112, 113). Therefore, “critical-looking” statistics (power laws, avalanches) are insufficient on their own (114, 99, 115, 116, 117) and broken symmetries underpin not all emergent behaviors (118, 119). Accordingly, strong claims of criticality, let alone emergence in living systems, require careful analysis (85, 120, 29, 92).
@@
Duality is an equivalence (often infrared-exact (121, 122)) between two seemingly different theories: a one-to-one “dictionary” that maps observables, states, and parameters so that all physical predictions coincide. Dualities reveal that theories with ontologically distinct variables—such as bosons versus fermions, or gravity versus no gravity—can describe the same underlying physics (121). Formally, a duality establishes a structural correspondence between the measurable quantities of two theories (a *-isomorphism) such that every observable, expectation value, and response function matches once external conditions—sometimes called “sources”—are translated accordingly. This agreement may accommodate local adjustments, such as contact terms or counterterms, reflecting different renormalization schemes (121).
In 1+11{+}1 dimensions, the bosonic sine–Gordon model is exactly dual to the fermionic Thirring model, exchanging solitonic waves for interacting particles and relating their couplings nonlinearly (123). In the Anti–de Sitter/conformal field theory correspondence, a conformal field theory without gravity on a $d$-dimensional boundary is holographically equivalent to a (d+1)(d{+}1)-dimensional bulk theory with gravity (48). Classic examples from statistical and gauge theories—such as high–low temperature mappings, electric–magnetic duality, and particle–vortex correspondence—further illustrate how distinct microscopic descriptions can encode the same macroscopic behavior (124, 125, 126, 127)
@@
Moving beyond physics, we consider how large-scale patterns emerge from local interactions in social and informational systems. Emergent phenomena possess causal power. In both the physical and social worlds, entities such as society and culture act as real forces shaping outcomes. Even in systems designed without formal rules—such as decentralized communities like Wikipedia or Stack Exchange—structures, hierarchies, and power distributions inevitably emerge, shaping the world in profound ways.